OpenStudy (karamarie02):

PLEASE ANSWER AS FAST AS POSSILE!!!!WILL FAN AND MEDAL!!!!!!!! what is 4/9 divided by 1/3?

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by the reciprocal

OpenStudy (anonymous):

\[\frac{4}{9}\div \frac{1}{3}=\frac{4}{9}\times \frac{3}{1}\] cancel first, multiply last

OpenStudy (karamarie02):

12/9? @satellite73

OpenStudy (anonymous):

well yes, but really you should cancel first and multiply last since you did not do that, you now have to reduce your fraction

OpenStudy (karamarie02):

@satellite73 so 4/3

jabez177 (jabez177):

Now you can't leave it as an improper fraction, you have to convert it to a mixed number. Divide 4 by 3. :)

OpenStudy (karamarie02):

@jabez177 so 1 1/3?

OpenStudy (anonymous):

yes \(\frac{4}{3}\)

jabez177 (jabez177):

Correct!

OpenStudy (karamarie02):

thanka